11. By solving an appropriate linear system, find the loop currents in the following circuit. 201 10v 100+ 3 T 14 В
Answer #1

20V 0 1+ 22 toop-1 ICOP-I 100P-II 100p-IEL 10v 11 44 45 }282 +140 452 22 B 3v 12c For loop-I by applying KVL we get 4(1,-12)=> 912 – 213 - 41, +1720 = -(41, +293 - 91, - 7) = 0 -) 41,-952+2 13 = 17 —-_>©for loop-fff by applying Kv2 equation we get -14 +463 +2 (13 - 12) = 0 => -212 +613 = 14 ->☺ we get From equation o 61, = 10 +412 f = 10 + 4 12 Now putting the value of I, in equ 4(104952) -9 12 +2 E3 = 17 => 3 + 12Iz in Ean we get putting the value of 2 + 623 = 14 => 6I3 =12 > Iz =2 Amp. o. The loop cuprents will be I = 1 Amp, 2. I2 =-1